Shift summary, June 19, 1999 Owl

Julie Whitmore (jaws@fnpxsr)
Sat, 19 Jun 1999 07:20:16 CDT

Shift Summary for owl shift, 06/19/99

On Shift:Mark Timko, Julie
Also present: Elliott, Elizabeth, Ed

Hours of beam delivered at all intensities: 0.25hr
Hours of beam delivered at nominal intensity: 0.25hr
Hours of E832 data (special runs not included): 0.25hr
Hours of special runs (laser scans, muon runs...): 0.25hr

Accounting of lost time: Access to debug and fix dead dynodes.
TeV F1 magnet quench.

========================================================================
Runs ended this shift:
Run 13661 - dynode debugging run
Run 13662 - Pedestal run.

Tapes brought to FCC this shift: None.

========================================================================
Accesses during shift:
Access to fix dynode channels 1066, 1216 and 392.

========================================================================
Overview of shift:

Majority of shift spent debugging and fixing dead dynodes in the calorimeter.
After access, beam returned for ~15min before a TeV injection quench at F1.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13661 00:15 Two inefficient HCC channels (1066, 1216) which also
show up as 3-5% low in gain drift plot. No signals
at ET monitoring. Ed suggested cycling HV.

00:40 Channel 1066 recovers, but not clear that
HV cycling cured it. Channel 1216 is dead at ET
monitoring.

01:30 Several attempts to fix 1216 by cycling HV fail.
Call for access. MCR sends beam to Meson because
we have exceeded 4hr limit of beam to our dump.

01:40 ECC and JW make access. No dynode signal for 1216 at
ET. Decide to try to debug dynode leaving all LV and
HV on. While trying to reseat dynode connector for
1216, JW blows fuse for DPMT board 1217. Swap DPMT
for channel 1217. Reseat dynode connector for 1216.

02:44 ECC and JW return. 1216 fixed. Now HCC bits for
1601, 1821, 1133 and 392. 1601 and 1821 have signals
at ET monitoring, 392 and 1133 are dead.

02:55 JW, ECC, SET go on access.

03:15 EB arrives. JW, ECC, SET fix 1133 by reseating dynode
connector. Channel 392 still dead.

03:35 EB swaps places with ECC on access. Swap cable
end of 392. Not fixed.

04:35 JW, EB, SET return from access. JW and EB return
to work on 392. Decide to patch dynode connector.
Finally fixed.

05:05 SET tweaks 6 Reg HVs (+/- 5V max change)

13662 05:45 Pedestal run after DPMT 1217 swap.

13663 05:50 Called ourselves out of access. MCR will retune beam
from Meson to KTeV. Beam returns at 5E12.

06:04 TeV injection quench @ F1. Est time for beam 30min.



========================================================================
Items to Follow up on Next Shift

Demacs program has hung on PLC. Need to figure out how to restart.